# Toward Reliable Sim-to-Real Predictability for MoE-based Robust Quadrupedal Locomotion

### v1

@@ III-B Mixture-of-Experts Representation Encoder
We substitute the student encoder in the CTS framework with the MoE network. This architecture comprises $K$ parallel expert subnetworks $\{E_{k}\}_{k=1}^{K}$ where each expert specializes in processing observation data under specific command types or environmental contexts. To coordinate these subnetworks, we incorporate a gating network $g$ that dynamically allocates weights $\omega_{k}$ based on the observation $\boldsymbol{o}_{t}$. These coefficients determine the relative contribution of each expert to the current state representation. Accordingly, the resulting latent state $\boldsymbol{z}_{s}$ of the student encoder is formulated as the weighted sum of all expert outputs:
$\boldsymbol{z}_{s}=\sum_{k=1}^{K}\omega_{k}E_{k}(\boldsymbol{o}_{t}),\quad\omega_{k}=\text{softmax}(g(\boldsymbol{o}_{t}))_{k}$ (2)
@@ III-C Reward Design
We utilize a consistent reward function structure for both the multi-terrain and the flat-ground high-speed locomotion models. The fundamental reward configurations are established based on established methodologies [24, 41, 50]. Building upon these foundations, we introduced a hip joint position reward to mitigate outward thigh abduction during rapid locomotion. Appendix Table VIII presents the comprehensive reward specifications. Within this framework, $\sigma$ denotes the velocity tracking precision parameter initialized to a value of 0.25. Additionally, the reward component $r^{\text{fr}}$ adopts the formulation from the CTS model [50] to incentivize adequate foot clearance during high-speed movement. For high-speed locomotion training on flat ground, we introduce an external hip symmetry reward $r^{\text{hs}}$ to regularize joint positions while executing longitudinal linear motion commands. This term ensures that the robot maintains symmetrical postures and is defined as follows:
@@ III-D Environment Configurations
We utilize the IsaacGym simulation environment [29] to train 8192 agents in parallel across diverse terrains. The experimental platform is the Unitree Go2 quadrupedal robot featuring 12 degrees of freedom. Motor PD control gains are specified as $k_{\text{p}}=20.0$ and $k_{\text{d}}=0.5$ for all joints. The system operates with a control frequency of 50Hz and a simulation frequency of 200Hz.
@@ IV The RoboGauge Predictive Assessment Framework
Figure 2: The RoboGauge evaluation architecture consists of three hierarchical stages. (A) Base Pipeline serves as a single evaluation environment by incorporating specific terrains and domain randomization. (B) Multi/Level Pipeline highlights the Multi/Level Pipeline where parallel evaluations across diverse random seeds. (C) Stress Pipeline depicts the Stress Pipeline which triggers comprehensive testing across the entire terrain suite to synthesize the final score.
@@ IV-C Hierarchical Scoring Methodology
We adopt a Worst-Case Mean aggregation strategy to evaluate performance across motion goals. This method involves averaging the lowest 50% of scores within each goal, effectively discounting high scores from non-challenging commands to concentrate the assessment on challenging maneuvers such as obstacle negotiation and gait transitions. Additionally, we compute the global mean and the average of the top 25% for broader reference as detailed in Appendix Table XII.
@@ V-A Metric Reliability of RoboGauge
We deployed the proposed model and baselines on a Unitree Go2 quadruped robot. We utilize a 12-camera Nokov Mars18H motion capture system operating at 90Hz to acquire real-time linear and angular velocity data across flat terrain and 10cm stairs by mounting five markers on the robot base. At the same time, we gather proprioceptive feedback and motor torques to derive the six specific metrics in Table II. To quantify the fidelity of these assessment methods, we compare the metric errors from both the training environment and our proposed framework against real-world ground truth. We specifically evaluate a model that exhibited high performance during training but suffered from significant sim-to-real degradation. As presented in Table III, the training environment consistently yields larger errors. Comprehensive scoring data provided in Table XI in the Appendix further confirms that errors obtained through our framework are markedly lower than those from standard training evaluations. These results demonstrate that our evaluation framework more accurately reflects real-world performance and provides a more dependable basis for model selection.
@@ V-B Comparison of Baselines under RoboGauge
Ours 0.67 $\textbf{0.66}\pm\textbf{0.24}$ $\textbf{0.78}\pm\textbf{0.25}$ $\textbf{0.77}\pm\textbf{0.25}$ 7.81
CTS $0.58$ $0.57\pm 0.23$ $0.70\pm 0.25$ $0.69\pm 0.25$ $6.81$
HIM $0.53$ $0.51\pm 0.26$ $0.63\pm 0.28$ $0.63\pm 0.28$ $5.94$
DWaQ $0.47$ $0.45\pm 0.27$ $0.57\pm 0.31$ $0.56\pm 0.30$ $5.15$
@@ VI Physical Deployment and Generalization
Figure 7: Robust locomotion during slope traversal and drop recovery. The left panel highlights a 1.7 s efficiency gain on $\mu=0.55$ slopes compared to the built-in RL baseline and the right frame verifies reliable recovery from 60 cm drops.
@@ VI-C Stability and Generalization
We validated the emergency recovery capabilities of the proposed model across two challenging real-world scenarios. First, the robot is subjected to external forces such as strong pushes or pulls where it shows great disturbance rejection by changing its center of mass and creating gaits to offset the impact. Fig. 9 and 17 show that the robot remains stable under continuous lateral pulls between 25 N and 40 N as well as sudden impulses of 85 N to 100 N where established baselines almost entirely fail to maintain balance. Second, when encountering a sudden loss of support the robot rapidly reconfigures its gait to secure its footing and prevent forward tumbling. Fig. 7 illustrates a successful recovery sequence from a 60 cm drop while Fig. LABEL:fig:unexpected_recovery depicts the natural transition to a stable posture after an unexpected fall from flat ground onto stairs.
@@ Appendix A RoboGauge Supplementary Material
A-A Implementation Details
@@
The BasePipeline orchestrates the interaction between the simulation engine sim, the evaluator gauge responsible for control commands and metric computation, and the locomotion model robot. Additionally, it manages exception handling, domain randomization, and the application of observation noise.
The MultiPipeline leverages multiprocessing to execute the BasePipeline across diverse seeds and domain randomization configurations while aggregating the output files. To determine the maximum navigable difficulty for a given terrain, the LevelPipeline identifies the highest level that the model traverses successfully across three separate random seeds.
In the quality score calculation Eq. 5, the metric weights are set to $w_{k}=2$ for task-completion metrics and $w_{k}=1$ to others. For the overlapping scoring function Eq. 6, the hyperparameters are set to $\alpha=0.09$ and $\beta=0.19$, which ensures the performance score is bounded within the range $[0,1]$.
@@ B-A Dynamic Velocity Tracking Precision Adjustment
The velocity commands $v$ pertain to longitudinal and lateral linear velocities as well as angular velocity commands. Table IX in the Appendix details the maximum velocity tracking coefficients $\sigma_{\text{max}}^{T_{i}}$ and the associated velocity adjustment ranges across diverse terrains.
@@ B-B Command Design
Direct training with the full command range of [1,1][-1,1] m/s across all terrains enables rapid progression through difficulty levels but frequently yields unstable gaits. Specifically, the robot often demonstrates erratic behaviors such as leaping and high-frequency leg motions. Conversely, training from low-speed commands facilitates the acquisition of stable locomotion patterns. We therefore introduce a command curriculum to address these issues, as detailed in Table X.
We observed that when the maximum command magnitude exceeds [1,1][-1,1] m/s, the robot fails to accurately track the target linear velocity on complex terrains such as wave, stairs, and obstacle. This tracking discrepancy induces instability during the training process. Therefore, we impose specific constraints on the maximum command range for individual terrains as detailed in Table IX. Notably, although these limits are strictly enforced during the training phase, no such restrictions are applied during hardware testing on the physical robot. Despite this discrepancy, the model follows commands that lie beyond the training distribution and demonstrates robust generalization capabilities.
@@ Appendix C Train configuration
TABLE VIII: Reward Function Specifications
@@
TABLE IX: Maximum Velocity Tracking Coefficients and Command Limits Across Terrains
@@
TABLE X: Command Curriculum Stages and Velocity Limits
@@ Appendix D Supplementary Experiment
TABLE XI: Comprehensive Evaluation: Real-World Measurements, Predicted Values, and Absolute Errors
@@
TABLE XII: RoboGauge detailed metrics for baselines
@@
Our 0.7633 0.7382 0.75 0.8036 0.7986 0.8009 0.7546 0.736 0.7435
CTS 0.6857 0.6604 0.6723 0.7237 0.7165 0.72 0.6777 0.6585 0.6666
HIM 0.6201 0.6008 0.6095 0.6541 0.6439 0.6493 0.6135 0.5975 0.6038
DreamWaQ 0.5611 0.5426 0.5512 0.5929 0.5835 0.5884 0.5508 0.5354 0.5415
TABLE XIII: RoboGauge detailed terrain scores for baselines
@@
Our 0.8102 0.7246 0.7578 0.7958 0.7103 0.7472 0.8737 0.7858 0.8219
CTS 0.8106 0.7309 0.7561 0.6185 0.5373 0.568 0.5785 0.4907 0.5289
HIM 0.4654 0.38 0.4125 0.7972 0.7153 0.7475 0.3646 0.2876 0.3227
DreamWaQ 0.3742 0.2991 0.3274 0.7419 0.6608 0.6916 0.206 0.1243 0.1645
Figure 14: Maximum terrain difficulty levels achieved by various models under a subset of friction coefficients (ranging from 0.5 to 2.5 in increments of 0.25).

### v2

@@ III-B Mixture-of-Experts Representation Encoder
We substitute the student encoder in the CTS framework with the MoE network. This architecture comprises $K$ parallel expert subnetworks $\{E_{k}\}_{k=1}^{K}$ where each expert specializes in processing observation data under specific command types or environmental contexts. To coordinate these subnetworks, we incorporate a gating network $g$ that dynamically allocates weights $\omega_{k}$ based on the observation sequence $\boldsymbol{o}_{t-H:t}=\left[\boldsymbol{o}_{t-H},\cdots,\boldsymbol{o}_{t}\right]^{T}$. These coefficients determine the relative contribution of each expert to the current state representation. Accordingly, the resulting latent state $\boldsymbol{z}_{s}$ of the student encoder is formulated as the weighted sum of all expert outputs:
$\boldsymbol{z}_{s}=\sum_{k=1}^{K}\omega_{k}E_{k}(\boldsymbol{o}_{t-H:t}),\quad\omega_{k}=\text{softmax}(g(\boldsymbol{o}_{t-H:t}))_{k}$ (2)
@@ III-C Reward Design
We utilize a consistent reward function structure for both the multi-terrain and the flat-ground high-speed locomotion models. The fundamental reward configurations are established based on established methodologies [24, 41, 50]. Building upon these foundations, we introduced a hip joint position reward to mitigate outward thigh abduction during rapid locomotion. Appendix Table IX presents the comprehensive reward specifications. Within this framework, $\sigma$ denotes the velocity tracking precision parameter initialized to a value of 0.25. Additionally, the reward component $r^{\text{fr}}$ adopts the formulation from the CTS model [50] to incentivize adequate foot clearance during high-speed movement. For high-speed locomotion training on flat ground, we introduce an external hip symmetry reward $r^{\text{hs}}$ to regularize joint positions while executing longitudinal linear motion commands. This term ensures that the robot maintains symmetrical postures and is defined as follows:
@@ III-D Environment Configurations
We utilize the IsaacGym simulation environment [29] to train 8192 agents in parallel across diverse terrains. The experimental platform is the Unitree Go2 quadrupedal robot featuring 12 degrees of freedom. Motor PD control gains are specified as $k_{\text{p}}=20.0$ and $k_{\text{d}}=0.5$ for all joints. The system operates with a control frequency of 50Hz and a simulation frequency of 200Hz. The length of the observation sequence $\boldsymbol{o}_{t-H:t}$ is set to 5 for MoE input. Algorithm configurations follow the CTS framework [50].
@@ IV The RoboGauge Predictive Assessment Framework
Figure 2: The RoboGauge evaluation architecture consists of three hierarchical stages. (A) Base Pipeline serves as a single evaluation environment by incorporating specific terrains and domain randomization. (B) Multi/Level Pipeline highlights the parallel evaluations across diverse random seeds. (C) Stress Pipeline triggers comprehensive testing across the entire terrain suite to synthesize the final score.
@@ IV-C Hierarchical Scoring Methodology
We adopt a Worst-Case Mean aggregation strategy to evaluate performance across motion goals. This method involves averaging the lowest 50% of scores within each goal, effectively discounting high scores from non-challenging commands to concentrate the assessment on challenging maneuvers such as obstacle negotiation and gait transitions. Additionally, we compute the global mean and the average of the top 25% for broader reference as detailed in Appendix Table XIII.
@@ V-A Metric Reliability of RoboGauge
We deployed the proposed model and baselines on a Unitree Go2 quadruped robot. We utilize a 12-camera NOKOV Mars18H motion capture system operating at 90Hz to acquire real-time linear and angular velocity data across flat terrain and 10cm stairs by mounting five markers on the robot base. At the same time, we gather proprioceptive feedback and motor torques to derive the six specific metrics in Table II. To quantify the fidelity of these assessment methods, we compare the metric errors from both the training environment and our proposed framework against real-world ground truth. We specifically evaluate a model that exhibited high performance during training but suffered from significant sim-to-real degradation. As presented in Table III, the training environment consistently yields larger errors. Comprehensive scoring data provided in Table XII in the Appendix further confirms that errors obtained through our framework are markedly lower than those from standard training evaluations. These results demonstrate that our evaluation framework more accurately reflects real-world performance and provides a more dependable basis for model selection.
@@ V-B Comparison of Baselines under RoboGauge
Ours 0.67 $\textbf{0.66}\pm\textbf{0.24}$ $\textbf{0.78}\pm\textbf{0.25}$ $\textbf{0.77}\pm\textbf{0.25}$ 7.85
CTS 0.58 $0.57\pm 0.24$ $0.70\pm 0.26$ $0.69\pm 0.26$ 6.83
HIM 0.54 $0.54\pm 0.25$ $0.64\pm 0.27$ $0.64\pm 0.27$ 6.19
DWaQ 0.50 $0.50\pm 0.26$ $0.61\pm 0.28$ $0.60\pm 0.28$ 5.74
@@ VI Physical Deployment and Generalization
Figure 7: Robust locomotion during slope traversal and drop recovery. The left panel highlights a 1.7 s efficiency gain on $\mu=0.71$ slopes compared to the built-in RL baseline and the right frame verifies reliable recovery from 60 cm drops.
@@ VI-C Stability and Generalization
We validated the emergency recovery capabilities of the proposed model across two challenging real-world scenarios. First, the robot is subjected to external forces such as strong pushes or pulls where it shows great disturbance rejection by changing its center of mass and creating gaits to offset the impact. Fig. 9 and 16 show that the robot remains stable under continuous lateral pulls between 25 N and 40 N as well as sudden impulses of 85 N to 100 N where established baselines almost entirely fail to maintain balance. Second, when encountering a sudden loss of support the robot rapidly reconfigures its gait to secure its footing and prevent forward tumbling. Fig. 7 illustrates a successful recovery sequence from a 60 cm drop while Fig. 17 depicts the natural transition to a stable posture after an unexpected fall from flat ground onto stairs.
@@ VII Conclusions and future work
Acknowledgements
The authors thank Tencent AI Arena and Unitree Robotics for providing the Go2 quadruped for the initial experiments. Thanks to Guangsheng Li for providing the training code for baseline DreamWaQ.
@@ Appendix A RoboGauge Supplementary Material
A-A Hyperparameter Configuration
In the quality score calculation Eq. 5, the metric weights are set to $w_{k}=2$ for task-completion metrics and $w_{k}=1$ to others. For the overlapping scoring function Eq. 6, the hyperparameters are set to $\alpha=0.09$ and $\beta=0.19$, which ensures the performance score is bounded within the range $[0,1]$. Domain randomizations include friction coefficients from $0.1$ to $1.0$ in increments of $0.1$. Terrain levels are designed with difficulty parameters $d$ ranging from $0.1$ to $1.0$ in increments of $0.1$ as detailed in Table LABEL:tab:terrain_design. The locomotion control objectives are configured as described in Table LABEL:table:goals.
A-B Implementation Details
@@
The BasePipeline (Fig. LABEL:fig:base_pipeline) orchestrates the interaction between the simulation engine sim, the evaluator gauge responsible for control commands and metric computation, and the locomotion model robot. Additionally, it manages exception handling, domain randomization, and the application of observation noise.
The MultiPipeline leverages multiprocessing to execute the BasePipeline across diverse seeds and domain randomization configurations while aggregating the output files. To determine the maximum navigable difficulty for a given terrain, the LevelPipeline (Fig. LABEL:fig:level_pipeline) identifies the highest level that the model traverses successfully across three separate random seeds.
@@ B-A Dynamic Velocity Tracking Precision Adjustment
The velocity commands $v$ pertain to longitudinal and lateral linear velocities as well as angular velocity commands. Table X in the Appendix details the maximum velocity tracking coefficients $\sigma_{\text{max}}^{T_{i}}$ and the associated velocity adjustment ranges across diverse terrains.
@@ B-B Command Design
Direct training with the full command range of [1,1][-1,1] m/s across all terrains enables rapid progression through difficulty levels but frequently yields unstable gaits. Specifically, the robot often demonstrates erratic behaviors such as leaping and high-frequency leg motions. Conversely, training from low-speed commands facilitates the acquisition of stable locomotion patterns. We therefore introduce a command curriculum to address these issues, as detailed in Table XI.
We observed that when the maximum command magnitude exceeds [1,1][-1,1] m/s, the robot fails to accurately track the target linear velocity on complex terrains such as wave, stairs, and obstacle. This tracking discrepancy induces instability during the training process. Therefore, we impose specific constraints on the maximum command range for individual terrains as detailed in Table X. Notably, although these limits are strictly enforced during the training phase, no such restrictions are applied during hardware testing on the physical robot. Despite this discrepancy, the model follows commands that lie beyond the training distribution and demonstrates robust generalization capabilities.
@@ Appendix C Train configuration
TABLE IX: Reward Function Specifications
@@
TABLE X: Maximum Velocity Tracking Coefficients and Command Limits Across Terrains
@@
TABLE XI: Command Curriculum Stages and Velocity Limits
@@ Appendix D Supplementary Experiment
TABLE XII: Comprehensive Evaluation: Real-World Measurements, Predicted Values, and Absolute Errors
@@
TABLE XIII: RoboGauge detailed metrics for baselines
@@
Our 0.7641 0.7359 0.7496 0.8056 0.8002 0.8028 0.7566 0.7371 0.7451
CTS 0.6850 0.6569 0.6708 0.7248 0.7161 0.7205 0.6781 0.6582 0.6666
HIM 0.6297 0.6012 0.6140 0.6686 0.6558 0.6629 0.6269 0.6080 0.6152
DreamWaQ 0.5976 0.5702 0.5829 0.6337 0.6178 0.6266 0.5913 0.5726 0.5798
TABLE XIV: RoboGauge detailed terrain scores for baselines
@@
Our 0.8268 0.7422 0.7744 0.7938 0.7104 0.7452 0.8847 0.7985 0.8349
CTS 0.8007 0.7210 0.7456 0.6682 0.5866 0.6179 0.5714 0.4830 0.5221
HIM 0.5624 0.4854 0.5149 0.7375 0.6638 0.6942 0.4627 0.3848 0.4197
DreamWaQ 0.5904 0.5148 0.5424 0.6839 0.6038 0.6358 0.3575 0.2808 0.3168
Figure 13: Maximum terrain difficulty levels achieved by various models under a subset of friction coefficients (ranging from 0.2 to 1.0 in increments of 0.1).
@@
Figure 17: The top panel shows the robot quickly adjusting its posture to safely descend when the ground ends at the edge. The middle plot depicts the contact force signals measured by the foot sensors. The bottom image illustrates the front foot height relative to the base calculated from forward kinematics. These results confirm the robustness of the policy and its capacity for adaptive gait transitions across diverse challenges.